# Compression fronts from fast radio bursts

### v1

@@ 9.2 Plasma compression and trapping by the FRB
where we used $dt=d\xi/(1-\beta_{\rm D})$ and $\kappa^{2}(1-\beta_{\rm D})=1+\beta_{\rm D}$. Note that this expression involves the FRB energy ${\cal E}$, assuming that the particle interacts with the entire radio packet, i.e. $t_{\rm cross}\leq r/c$, which corresponds to $C\leq C_{\rm trap}$. [^8] A “deep” trapping regime $t_{\rm cross}\gg r/c$ is possible if the plasma is trapped in a small leading part of the wave packet, at $\xi_{\rm trap}\ll T$. In this case, ${\cal E}$ in Equation (116) should be replaced by the part of the FRB energy ${\cal E}_{\rm trap}$ contained in the leading part $0<\xi<\xi_{\rm trap}$. Self-consistent $\xi_{\rm trap}$ and ${\cal E}_{\rm trap}$ depend on the packet shape.
Using Equations (115) and (116) in Equation (114), we find
$C_{0}^{1/2}\sim\frac{\tilde{\gamma}}{\sigma_{\rm u}}\left(\frac{3}{2}+\psi\right),\qquad\psi\equiv\sqrt{\frac{\sigma_{\rm T}{\cal E}\sigma_{\rm u}}{4\pi r^{2}mc^{2}\kappa_{\rm u}}}.$ (117)
Here, $\tilde{\gamma}=\min\{\tilde{\gamma}_{\rm stoch},\tilde{\gamma}_{\rm rad}\}$ depends on $C=C_{0}$ (section 9.1), and one can find the consistent solution for $\tilde{\gamma}$ and $C_{0}$:
$\tilde{\gamma}=\left\{\begin{array}[]{lr}\tilde{\gamma}_{\rm stoch}&\quad\psi<0.7\\ \tilde{\gamma}_{\rm rad}&\quad\psi>0.7\end{array}\right.$ (118)
$C_{0}\sim\left[\frac{c\kappa_{\rm u}}{r_{e}\omega}\right]^{6/5}\!\!\frac{\alpha^{2/5}}{16\sigma_{\rm u}^{4}}\!\times\!\left\{\begin{array}[]{lr}(3\psi)^{12/5}(3+2\psi)^{14/5}&\psi<0.7\\ (3+2\psi)^{4}&\psi>0.7\end{array}\right.$ (119)
with the parameter $\alpha$ given by Equation (88).
The obtained solution holds if it gives $C_{0}\gg 1$, and this condition is satisfied at all $r<r_{\rm stoch}$. Furthermore, $C_{0}$ exceeds $C_{\rm trap}$. Indeed, consider the smallest $C_{0}^{1/2}\sim 3\tilde{\gamma}_{\rm stoch}/2\sigma_{\rm u}$ (Equation 117), which is found in the non-radiative regime $\psi\ll 0.7$ (small ${\cal E}$). Then, using Equation (110) for $\tilde{\gamma}_{\rm stoch}$, we find
$C_{0}\sim\frac{10^{6}\,{\cal E}_{39}^{6/5}L_{\rm w,37}^{1/5}}{\nu_{9}^{8/5}\eta_{4}^{14/5}R_{\rm LC,10}^{14/5}}\!\!\times\!\left\{\begin{array}[]{lr}1&r<R_{\rm sat}\\ (R_{\rm sat}/r)^{14/5}&r>R_{\rm sat}\end{array}\right.$ (120)
where $R_{\rm sat}=\eta^{1/3}R_{\rm LC}\approx 2\times 10^{11}\,\eta_{4}^{1/3}R_{\rm LC,10}$ cm is the radius where the wind Lorentz factor $\gamma_{\rm u}$ approaches $\eta^{1/3}$ (Equation 85). It is easy to verify that $C_{0}>C_{\rm trap}$ at all $r<r_{\rm stoch}$. So, FRBs emitted in the wind at $r<r_{\rm stoch}$ immediately trap the wind plasma.
Two aspects of the trapping regime should be noted:
(1) The large value of $C_{0}$ is found using the full FRB energy ${\cal E}\sim LT$ in the parameter $\psi$. When $C_{0}\gg C_{\rm trap}$, the plasma becomes trapped in a small leading part of the radio wave packet with ${\cal E}_{\rm trap}\ll{\cal E}$ (see footnote 8), avoiding interaction with its main part.
(2) The trapping regime holds unless the FRB becomes strongly damped by its interaction with the wind. Damping allows the plasma to move through the wave packet while consuming its energy, avoiding trapping. The parameter space where FRBs are damped is evaluated below.
10 FRB energy losses
Consider first FRB energy losses $\Delta{\cal E}^{\prime}$ in frame ${\cal K}^{\prime}$ where the upstream plasma is at rest. The corresponding momentum losses are $\Delta{\cal E}^{\prime}/c$.
Let the radio wave packet interact with $\Delta{\cal N}$ particles. The momentum lost by the radio wave is the sum of momentum gained by the MHD flow and momentum waisted by the energized plasma particles to radiation,
$\frac{\Delta{\cal E}^{\prime}}{c}=\Delta{\cal N}\,\frac{T^{t^{\prime}r^{\prime}}_{\rm MHD}}{cn^{\prime}}+\beta_{\rm D}^{\prime}\frac{{\cal E}_{\rm rad}^{\prime}}{c}.$ (121)
The radiated energy per particle is $\epsilon_{\rm rad}^{\prime}mc^{2}\equiv{\cal E}_{\rm rad}^{\prime}/\Delta{\cal N}$. This gives
$\frac{\Delta{\cal E}^{\prime}}{mc^{2}\Delta{\cal N}}\approx\beta_{\rm D}^{\prime}\left[\gamma_{\rm D}^{\prime}(1+w+\sigma_{\rm u}\kappa^{\prime})+\epsilon_{\rm rad}^{\prime}\right].$ (122)
Note that $\beta_{\rm D}^{\prime}=(\kappa^{\prime 2}-1)/(\kappa^{\prime 2}+1)$ and $\kappa^{\prime}\approx\tilde{C}$, where $\tilde{C}=\tilde{\rho}/\tilde{\rho}_{\rm u}$ is the proper compression factor. Thus, Equation (122) yields
$\frac{\Delta{\cal E}^{\prime}}{mc^{2}\Delta{\cal N}}\approx(\tilde{C}^{2}-1)\left(\frac{1+w}{2\tilde{C}}+\frac{\sigma_{\rm u}}{2}+\frac{\epsilon_{\rm rad}^{\prime}}{\tilde{C}^{2}+1}\right).$ (123)
Proper compression $\tilde{C}$ is related to compression $C$ in the lab frame by $C\approx\tilde{C}^{2}$ (Equation 98). Wave energy in the lab frame is ${\cal E}=\kappa_{\rm u}{\cal E}^{\prime}$, so $\Delta{\cal E}^{\prime}=\Delta{\cal E}/\kappa_{\rm u}$. Radiative losses in the lab frame and frame ${\cal K}^{\prime}$ are also related by ${\cal E}_{\rm rad}/{\cal E}_{\rm rad}^{\prime}\approx\kappa_{\rm u}$. Using these relations, one can rewrite Equation (123) in terms of the lab-frame quantities:
$\frac{\Delta{\cal E}}{mc^{2}\Delta{\cal N}\kappa_{\rm u}}\approx(C-1)\left[\frac{1+w}{2\sqrt{C}}+\frac{\sigma_{\rm u}}{2}+\frac{\epsilon_{\rm rad}}{(C+1)\kappa_{\rm u}}\right].$ (124)
@@ 10 FRB energy losses
Equations (124) and (125) determine the FRB energy losses $\Delta{\cal E}$ anywhere in the wind, expressed as a function of the compression factor $C$. Enthalpy $w\approx(3/2)\tilde{\gamma}$ is controlled by the heating regime: $\tilde{\gamma}=(1+a^{2})^{1/2}$ at $r>r_{\rm stoch}$, and $\tilde{\gamma}$ is given by Equation (110) at $r<r_{\rm stoch}$.
It is easy to verify that losses are negligible, $\Delta{\cal E}\ll{\cal E}$, in the zone of regular heating $r>r_{\rm stoch}$. Thus, FRB damping in the wind may only occur in the zone of stochastic heating, $r<r_{\rm stoch}$.
Using $w\approx(3/2)\tilde{\gamma}\gg 1$ and $C\gg 1$, we find from Equation (124):
$\Delta{\cal E}\approx\frac{rL_{\rm w}}{2c\,\eta\gamma_{\rm u}}\left[\frac{3}{2}\tilde{\gamma}\sqrt{C}+\sigma_{\rm u}C+\frac{\sigma_{\rm T}\tilde{\gamma}^{2}{\cal E}}{4\pi r^{2}mc^{2}\gamma_{\rm u}}\right].$ (126)
For an FRB with luminosity $L$ and energy ${\cal E}\sim LT$, we find $\tilde{\gamma}(r,C)$ from Equation (110) at radii $r<r_{\rm stoch}$, and then find $\Delta{\cal E}(r,C)$ from Equation (126).
The FRB is strongly damped if $\Delta{\cal E}$ approaches ${\cal E}$. Damping begins in the leading part of the FRB $\xi\ll T$, as the plasma entering the radio wave packet consumes its local energy ${\cal E}(\xi)\sim 4\pi r^{2}U(\xi)\,c\,\xi$ and makes the deep trapping regime $C\gg C_{\rm trap}$ unsustainable. Damping succeeds in killing the FRB when it enables the plasma to keep going in $\xi$ and crossing the entire wave packet, which corresponds to $\Delta{\cal E}\sim{\cal E}$ and marginal trapping, $C\approx C_{\rm trap}$. Therefore, the condition for marginal damping is
@@
It defines the boundary of the parameter space where FRBs are strongly damped. One can show that $\Delta{\cal E}(r,C_{\rm trap})/{\cal E}$ is a decreasing function of ${\cal E}$, so strong damping occurs for FRBs with energies ${\cal E}$ below a critical value ${\cal E}_{\rm esc}(r)$ defined by Equation (127) and the corresponding critical luminosity is $L_{\rm esc}\sim{\cal E}_{\rm esc}/T$. The damping region on the $r$-$L$ plane is shown in Figure 8.
Figure 8: Regimes of FRB propagation through the wind ($r>R_{\rm LC}$), shown on the $r$-$L$ plane, for FRBs with emission frequency $\nu=1$ GHz and duration $T=1$ ms. The two plots assume the wind energy parameter $\eta=10^{4}$ (left) and $\eta=10^{3}$ (right). The magenta region $r<R_{\rm LC}$ indicates the closed magnetosphere. FRBs emitted inside the magnetosphere experience strong radiative damping (Beloborodov 2024). FRBs emitted at radii $R_{\rm LC}<r<r_{\rm stoch}$ induce fast stochastic heating and compression of the wind. In the light purple region, FRBs trap the wind plasma, carrying it in a leading part of the wave packet and avoiding significant energy losses. In the deep purple region, FRBs are damped: their energy is consumed by heating and compressing the wind plasma, with radiative losses. The colored dots indicate the dominant damping channel: radiative losses ${\cal E}_{\rm rad}$ (red dots), energy kept by the plasma ${\cal E}_{\rm p}$ (cyan dots) and energy spent to compress the wind magnetic field ${\cal E}_{\rm mag}$ (yellow dots). In the olive zone, $r_{\rm stoch}<r<r_{1}$, the wind is still heated to a relativistic enthalpy $w\gg 1$; however, here heating occurs in the regular regime (particles oscillate in the radio wave instead of stochastically gaining energy) with the moderate $w\approx(3/2)a$ and a weaker compression. At radii $r_{\star}<r<r_{1}$, the compression factor is $C=C_{\star}\approx 1+a_{0}^{2}$. The FRB strength parameter $a=a_{0}/\sqrt{2}$ drops below unity outside $r_{1}$ (grey region).
@@
where ${\cal E}_{\rm p}\propto\tilde{\gamma}$ is the energy kept by the plasma particles, ${\cal E}_{\rm mag}\propto\sigma_{\rm u}$ is the energy used to compress the wind magnetic field, and ${\cal E}_{\rm rad}\propto\tilde{\gamma}^{2}$ is the radiative losses. Let us compare radiative losses ${\cal E}_{\rm rad}$ with ${\cal E}_{\rm p}$ and ${\cal E}_{\rm mag}$.
@@
It is easy to see that ${\cal E}_{\rm rad}>{\cal E}_{\rm p}$ if $\tilde{\gamma}_{\rm rad}<\tilde{\gamma}_{\rm stoch}$; in this regime, the plasma radiates more energy than it keeps. This occurs at radii that satisfy the condition given in Equation (111). With $C=C_{\rm trap}$, it can be stated as
${\cal E}_{\rm rad}>{\cal E}_{\rm p}:\quad r<2\times 10^{11}\,\frac{{\cal E}_{39}^{20/47}T_{\rm ms}^{5/47}L_{\rm w,37}^{1/47}}{\gamma_{\rm u}^{4/47}\nu_{9}^{8/47}}\,{\rm cm}.$ (129)
The ratio ${\cal E}_{\rm rad}/{\cal E}_{\rm mag}$ equals the ratio of the third and second terms in the brackets in Equation (126). Using $C=C_{\rm trap}$ and $\tilde{\gamma}=\tilde{\gamma}_{\rm rad}$ (i.e. assuming ${\cal E}_{\rm rad}>{\cal E}_{\rm p}$), we find
${\cal E}_{\rm rad}>{\cal E}_{\rm mag}:\quad r<10^{11}\gamma_{\rm u}^{16/27}\frac{{\cal E}_{39}^{10/27}T_{\rm ms}^{5/27}L_{\rm w,37}^{1/27}}{\eta_{4}^{10/27}\nu_{9}^{8/27}}\,{\rm cm}.$ (130)
${\cal E}_{\rm rad}$ dominates the FRB losses at radii that satisfy both conditions stated in Equations (129) and (130).
11 Summary
The basic physics problem studied in this paper is how a strong ($a_{0}\gg 1$) electromagnetic wave packet accelerates an MHD fluid. For weak electromagnetic waves ($a_{0}\ll 1$), the acceleration effect is well known as the radiation pressure force, used in astrophysics e.g. to define the Eddington limit. For strong waves, the acceleration effect was previously calculated in unmagnetized plasmas; the $e^{\pm}$ plasma then becomes compressed in the wave packet by the factor $C\approx 1+a_{0}^{2}$. A similar problem for a magnetized plasma was poorly understood. The present paper developed a method to solve this problem, obtained sample numerical solutions, and provided a physical picture of the wave-plasma interaction. Two key facts help solve the problem: (1) the wave strength parameter $a_{0}$ controls the plasma enthalpy as shown in section 3, and (2) quantity ${\cal Q}_{-}={\cal E}-c{\cal P}$ (where ${\cal E}$ is energy and ${\cal P}$ is momentum) is conserved in the wave-plasma interaction because the wave has ${\cal Q}_{-}=0$ and cannot transfer $\Delta{\cal Q}_{-}\neq 0$ to the MHD fluid.
A first estimate would suggest that plasma with a sufficiently large magnetization $\sigma$ is not strongly compressed by the electromagnetic wave, since the deposited momentum is proportional to the small plasma density while compression requires work to be done against the large background magnetic pressure (Beloborodov 2021; see also section 6.1). The actual solution demonstrates that this expectation holds initially, when the wave packet begins to interact with the plasma. Later, the propagating wave gradually develops a large compression factor $C$ due to a curious “memory” effect described in section 6. Given a sufficient time $t_{\star}\propto C\sigma$ the compression front relaxes to a steady state with $C=C_{\star}\approx 1+a_{0}^{2}$ independent of $\sigma$. However, large $C_{\star}$ (implied by large $a_{0}$) is not reached if $t_{\star}$ exceeds the wave propagation time.
Besides the peculiar gradual evolution of the propagating wave packet, the background magnetic field brings another special feature that dramatically changes the wave-plasma interaction: compression increases the gyro-frequency and triggers strong stochastic heating of the plasma by the electromagnetic wave. The heating boosts momentum deposition, enhancing acceleration and compression of the MHD fluid in the wave packet.
We have used the obtained results to examine FRB interaction with a magnetized wind flowing from the central object (a neutron star). Our conclusions are summarized below and illustrated in Figure 8.
FRBs propagating at radii $R_{\rm LC}<r<r_{\rm stoch}$ have a huge effect on the wind medium. They heat the wind up to an average Lorentz factor $\tilde{\gamma}\sim 10^{3}a_{0}$ (measured in the fluid rest frame), and accelerate it to a bulk Lorentz factor exceeding $\gamma_{\rm trap}\sim(r/cT)^{1/2}$, which implies trapping of the wind plasma inside the radio wave packet. The energy per particle in the heated plasma $\gamma_{\rm trap}\tilde{\gamma}mc^{2}$ (measured in the static lab frame) is in the TeV range. Thus, the FRB-wind interaction converts the radio-wave energy into high energy particles, which experience radiative losses. A fraction of FRB energy is also spent to compress the wind magnetic field. Only sufficiently bright FRBs survive the energy losses (Figure 8).
The FRB-wind interaction changes at larger radii: at $r=r_{\rm stoch}$ (Equation 104) there is a sharp transition from stochastic heating to regular heating (particle oscillation in the wave), which sustains a moderate $\tilde{\gamma}\sim a_{0}$. Energy losses of the radio wave are negligible at $r>r_{\rm stoch}$. FRBs emitted in this zone still induce strong bulk acceleration of the wind plasma. Here, instead of trapping the plasma, the FRB drives an MHD compression wave in the wind. At $r>r_{\star}$ (Equation 95), the compression wave is quasisteady ($r/c\gg t_{\star}$), and the compression factor equals $C_{\star}\approx 1+a_{0}^{2}$.
@@ 11 Summary
The results summarized in Figure 8 imply interesting constraints on FRB emission models. In particular, they exclude scenarios that place the radio source at radii $r<r_{\rm damp}\sim 10^{11}$ cm and rely on free FRB propagation through the magnetar wind. Thus, the previously found damping zone $r<R_{\rm LC}$ (Beloborodov, 2021, 2024; Golbraikh & Lyubarsky, 2023) is extended to $r_{\rm damp}\sim 10^{11}$ cm. The exact value of $r_{\rm damp}$ varies around $10^{11}$ cm for the relevant values of wind parameter $\eta=L_{\rm w}/\dot{\cal N}mc^{2}$ around $10^{4}$. We find that $r_{\rm damp}$ slowly decreases with increasing $\eta>10^{4}$. The compression and damping effects are particularly strong in winds with $\eta<10^{4}$.

### v2

@@ X FRB energy losses
$\Delta{\cal E}\approx\Delta{\cal N}\,\frac{T^{tt}_{\rm MHD}}{n}+\Delta{\cal N}{\cal E}_{e},$ (114)
where ${\cal E}_{e}$ is the radiated energy per particle. In this relation, we have neglected the initial energy of the wind compared to its gain from the FRB. Using $T_{\rm MHD}^{tt}=\tilde{\rho}c^{2}h\gamma_{\rm D}^{2}-P$ and dropping the term $P$ (it is smaller than the main term by the factor $\sim\gamma_{\rm D}^{-2}\ll 1$) one finds
$\frac{T^{tt}_{\rm MHD}}{n}\approx mc^{2}\gamma_{\rm D}(1+w+\sigma).$ (115)
Radiative losses of a particle interacting with the radio wave packet are derived in Appendix D in the regimes of regular oscillation and stochastic heating. For the MHD flow with $\gamma_{\rm D}\gg 1$ and $\tilde{\gamma}\gg 1$ they simplify to
$\frac{\Delta{\cal E}_{e}}{\Delta t}\approx\frac{c\sigma_{\rm T}U\tilde{\gamma}^{2}}{2\kappa^{2}}.$ (116)
Then, we find
${\cal E}_{e}=\!\int\!\frac{\Delta{\cal E}_{e}}{\Delta t}\,dt\approx\!\int\!\frac{c\sigma_{\rm T}\tilde{\gamma}^{2}Ud\xi}{4}\approx\frac{\sigma_{\rm T}\tilde{\gamma}^{2}{\cal E}}{16\pi r^{2}},$ (117)
[…]
Using $\gamma_{\rm D}/\gamma_{\rm u}\approx\sqrt{C}\gg 1$ and $\sigma/\sigma_{\rm u}=\tilde{\rho}/\tilde{\rho}_{\rm u}\approx\sqrt{C}$ (Equation 97), we obtain
$\frac{\Delta{\cal E}}{\Delta{\cal N}mc^{2}\gamma_{\rm u}}\approx\sqrt{C}\,w+C\sigma_{\rm u}+\frac{\sigma_{\rm T}{\cal E}\,\tilde{\gamma}^{2}}{16\pi r^{2}mc^{2}\gamma_{\rm u}}.$ (118)
@@
Equations (118) and (119) determine the FRB energy loss $\Delta{\cal E}$, expressed as a function of the compression factor $C$. It is easy to verify that losses are negligible, $\Delta{\cal E}\ll{\cal E}$, in the zone of regular heating $r>r_{\rm stoch}$. Thus, FRB damping in the wind may only occur in the zone of stochastic heating, $r<r_{\rm stoch}$. Then, $w\approx(3/2)\tilde{\gamma}$ and Equation (118) gives
$\Delta{\cal E}\approx\frac{rL_{\rm w}}{2c\,\eta\gamma_{\rm u}}\left[\frac{3}{2}\tilde{\gamma}\sqrt{C}+\sigma_{\rm u}C+\frac{\sigma_{\rm T}\tilde{\gamma}^{2}{\cal E}}{16\pi r^{2}mc^{2}\gamma_{\rm u}}\right].$ (120)
[…]
Figure 10: Regimes of FRB propagation through the wind ($r>R_{\rm LC}$), shown on the $r$-$L$ plane, for FRBs with frequency $\nu=1$ GHz and duration $T=1$ ms. The two plots assume the wind energy parameter $\eta=10^{4}$ (left) and $\eta=10^{3}$ (right). Magenta region $r<R_{\rm LC}$ indicates the closed magnetosphere; FRBs emitted in this region are known to experience strong radiative damping (Beloborodov, 2024). FRBs are also damped in the wind zone $r>R_{\rm LC}$ studied in the present paper, in particular in the deep purple region. Their energy is consumed by the stochastic heating and compression of the wind plasma, with radiative losses. Colored dots indicate the dominant damping channel: radiative losses ${\cal E}_{\rm rad}$ (red dots), energy kept by the plasma ${\cal E}_{\rm p}$ (cyan dots), and energy spent to compress the wind magnetic field ${\cal E}_{\rm mag}$ (yellow dots). Stochastic heating occurs at $r<r_{\rm stoch}$. In the light purple region, FRBs trap the wind plasma, carrying it in a leading part of the wave packet and avoiding significant energy losses. In the olive zone, $r_{\rm stoch}<r<r_{1}$, FRBs induce regular particle oscillations with $\tilde{\gamma}\gg 1$ and drive a compression wave through the wind, with a decreasing compression factor $C$. At radii $r>r_{\star}$, the compression wave is locally quasisteady and $C=C_{\star}\approx 1+a^{2}$. The FRB strength parameter $a$ drops below unity outside $r_{1}$ (grey region).
[…]
@@
[…]
@@
where ${\cal E}_{\rm p}\propto\tilde{\gamma}$ is the energy kept by the plasma particles, ${\cal E}_{\rm mag}\propto\sigma_{\rm u}$ is the energy used to compress the wind magnetic field, and ${\cal E}_{\rm rad}\propto\tilde{\gamma}^{2}$ is the radiative losses. The largest of these three contributions defines the main damping channel for the FRB; it is indicated in Figure 10. It is easy to see that ${\cal E}_{\rm rad}>{\cal E}_{\rm p}$ if $\tilde{\gamma}_{\rm rad}<\tilde{\gamma}_{\rm stoch}$; in this regime, the plasma radiates more energy than it keeps. This occurs at radii that satisfy the approximate condition stated in Equation (111) with $C=C_{\rm trap}$.
XI Summary
@@
A first estimate would suggest that plasma with a sufficiently large magnetization $\sigma$ is not strongly compressed by the radio wave, since the deposited momentum is proportional to the small plasma density while compression requires work to be done against the large background magnetic pressure (Beloborodov 2021; see also section VI). The actual solution demonstrates that this expectation holds initially, when the wave packet begins to interact with the plasma. If the propagating radio wave is given sufficient time, it gradually develops a much larger compression factor $C$. After time $t_{\star}\propto C\sigma$ the compression front relaxes to a steady state with $C=C_{\star}\approx 1+a^{2}$, which is independent of $\sigma$.
Besides the peculiar gradual evolution of the propagating compression front, the background magnetic field brings another special feature: it enables powerful stochastic heating of the plasma by the radio wave, in particular when compression increases the local gyro-frequency. The heating boosts momentum deposition, further enhancing acceleration and compression of the MHD fluid in the wave packet.
We have used the obtained results to examine FRB interaction with a magnetized wind flowing from the neutron star. Our conclusions are summarized below and illustrated in Figure 10.
@@
The FRB-wind interaction changes at larger radii: at $r=r_{\rm stoch}\sim 10^{12}$ cm (Equation 105) there is a sharp transition from stochastic heating to regular particle oscillations in the radio wave, with a moderate Lorentz factor $\tilde{\gamma}\sim a$. FRB energy losses are negligible at $r>r_{\rm stoch}$. FRBs emitted in this zone still induce strong bulk acceleration of the wind plasma, but no longer trap it, and continue to drive an MHD compression wave in the wind. The compression wave becomes locally quasisteady at $r>r_{\star}\sim 10^{13}L_{42}^{1/3}$ cm (Equation 93), with compression factor following the simple expression $C_{\star}\approx 1+a^{2}$. In this region, our results agree with Sobacchi et al. (2024).
It should be emphasized that we find the compression factor $C\ll C_{\star}$ at $r\ll r_{\star}$, where the timescale for the compression front relaxation toward $C_{\star}$ exceeds the FRB propagation time $r/c$. Therefore, our results do not support the suggestion of Sobacchi et al. (2024) that FRBs emitted at $r\lesssim 10^{12}$ cm lose energy to the huge compression $C_{\star}$ of the magnetized wind. [^14] Note also that energy consumption implied by $C_{\star}\approx 1+a^{2}$ is reversible: when the MHD fluid exits behind the wave packet (where $C_{\star}$ comes back to unity) it returns energy to the FRB.
Our results still imply a significant constraint on the FRB emission radius (Figure 10). It is caused by the enormous stochastic heating of the wind exposed to the FRB at small $r$, which leads to catastrophic energy losses of the FRB. As a result, the previously found damping zone $r<R_{\rm LC}$ (Beloborodov, 2021, 2024; Golbraikh & Lyubarsky, 2023) is now extended to $r_{\rm damp}\sim 10^{11}$ cm. This will challenge any FRB scenarios that place radio emission at $r<10^{11}$ cm and rely on its escape through the magnetar wind. The exact $r_{\rm damp}$ depends on the wind parameter $\eta=L_{\rm w}/\dot{\cal N}mc^{2}$, which may vary around $\eta\sim 10^{4}$. We find that $r_{\rm damp}$ slowly decreases with increasing $\eta>10^{4}$. FRB damping is particularly strong in winds with $\eta<10^{4}$.
[…]